OpenStudy (anonymous):

when the cost of a car is multiplied by 0.07 the result is 1720 how much is the car

OpenStudy (anonymous):

If a*b=c, how do you get a? Solve for a, so a=c/b a= cost of the car, b=0.07 and c=1720